the Fahrenheit temperature, F, is found using formula F = 9/5C+32, where C the temperature in Celsius, determine the formula for C and justify each step.
Answer by nerdybill(7384)


Start with:
F = 9/5C+32
subtract 32 from both sides:
F - 32 = 9/5C
multiply both sides by 5:
5F - 160 = 9C
divide both sides by 9:
(5F - 160)/9 = C